Results from MathSciNet: Mathematical Reviews on the Web \begin{matrix}c \\ ◯ \end{matrix} Copyright American Mathematical Society 2007

MR0529598 (80h:32041) 32G13 Forster, Otto; Knorr, Knut

⋆Konstruktion verseller Familien kompakter komplexer Räume. (German) [Construction of versal families of compact complex spaces] Lecture Notes in Mathematics, 705. Springer, Berlin, 1979. vii+141 pp. $9.00. ISBN 3-540-09122-* Die Verfasser geben einen neuen Beweis für das folgende Theorem: Zu jedem kompakten komplexen Raum X existiert eine verselle Deformation. Dieser Satz wurde zuerst 1974 von Douady und Grauert mit unterschiedlichen Methoden bewiesen; 1976 publizierte Palamodov ebenfalls einen Beweis. Angesichts der grossen Bedeutung dieses Theorems ist es zu begrüssen, dass nun eine weitere Lösung des Modulproblems für kompakte komplexe Räume gegeben wird. Während Douady und Grauert die Theorie der Banach-analytischen Räume verwenden, steht in dieser Arbeit die Potenzreihenmethode im Mittelpunkt. Es sei X ein kompakter komplexer Raum und S ein komplexer Raumkeim mit ausgezeichnetem Punkt s_{0}. Eine platte Familie komplexer Räume π: Y → X mit π^{−1}(s_{0}) = X heisst eine Deformation von X. Die Deformation heisst vollständig, wenn es zu jeder weiteren Deformation Y′ → S′ von X einen Morphismus

\begin{matrix}Y′ & ⟶^{f} & Y \\ ↓ & & ↓ \\ S′ & ⟶^{φ} & S \end{matrix}

gibt. Sie heisst versell, wenn ausserdem das Differential dφ im ausgezeichneten Punkt eindeutig bestimmt ist. Schlessinger, H. W. Schuster und Wavrik haben gezeigt: Zu jedem kompakten komplexen Raum existiert eine formale Deformation, die formal versell ist. Eine (konvergente) Deformation ist genau dann versell, wenn sie formal versell ist.—In der vorliegenden Arbeit wird von der formalen versellen Deformation ausgegangen und gezeigt, dass die Erweiterung um jeweils eine Ordnung so durchgeführt werden kann, dass man eine konvergente Deformation erhält.

Dazu wird zunächst eine Hindernistheorie entwickelt. Zu X wählt man eine endliche offene Überdeckung {X_{i}} und Einbettungen von X_{i} in Polyzylinder D_{i} ⊂ C^{n}. Um die Plattheit kontrollieren zu können, wählt man ausserdem eine freie Auflösung der Strukturgarbe von X_{i}. Zusammen mit Verheftungsabbildungen und Komplexisomorphismen über den Durchschnitten liefert dies die sogenannte analytische Einspannung von X. Dadurch kann man die Verheftung der lokalen Deformationen beschreiben. Die globalen Deformationen werden durch nicht-abelsche Cozyklen dargestellt. Die analytische Einspannung liefert einen Extensions-komplex E^{·}; es wird gezeigt, dass H^{1}(E^{·}) endlich-dimensional und isomorph zum Vektorraum der infinitesimalen Deformationen von X ist. Die Hindernisse gegen die Fortsetzung einer Deformation liegen in H^{2}(E^{·}). Die von Schlessinger angegebene formale verselle Deformation wird nun bei Fortsetzung zur nächsthöheren Ordnung so abgeändert, dass man eine konvergente Deformation erhält. Um die Hindernisse abschätzen zu können, geht man zu feineren Überdeckungen über. Damit die Fortsetzung unbegrenzt oft durchgeführt und abgeschätzt werden kann, ist es erforderlich, den die Deformation darstellenden Cozyklus auf der feineren Überdeckung durch einen auf einer gröberen Überdeckung zu ersetzen. Dazu sind komplizierte Glättungstechniken erforderlich. Es wird eine Differentialrechnung in induktiv normierten Räumen entwickelt. Eine wesentliche Rolle spielen die von Grauert hergeleiteten Aussagen über reduzierende Systeme und Weierstrass-Familien. Die Verfasser beweisen ein Aufspaltungslemma, ein Theorem B für Deformationen, ein Reduktions- und Überschusslemma und Glättungssätze. Damit wird dann der Konvergenzbeweis geführt.

H. Kerner (Bayreuth)

MR0460703 (57 696) 32C15 Forster, O.; Knorr, K. Espaces analytiques relatifs. (French) Fonctions analytiques de plusieurs variables et analyse complexe (Colloq. Internat. CNRS, No. 208, Paris 1972), pp. 51–52, ‘‘Agora Mathematica’’, No. 1, Gauthier-Villars, Paris, 1974. From the authors’ text: “Let but de l’exposé est de présenter quelques techniques utilisées pour étudier les espaces analytiques relatifs. Pour les détails nous renvoyons au Chapitre I de l’article des auteurs [Invent. Math. 16 (1972), 113–160; MR0325998 (48 #4344)].”

{For the entire collection see MR0450592 (56 #8885).} MR0374495 (51 10695) 32G05 Forster, O.; Knorr, K. Über die Deformationen von Vektorraumbündeln auf kompakten komplexen Räumen. (German) Math. Ann. 209 (1974), 291–346. The authors prove that every holomorphic vector bundle over a compact analytic space possesses a versal deformation. The existence of a formal versal deformation is used (though not proved; the authors refer to two previous papers [M. Schlessinger, Trans. Amer. Math. Soc. 130 (1968), 208–222; MR0217093 (36 #184); H. W. Schuster, “Formale Deformationstheorien”, Habilitationsschrift, Ludwig-Maximilian Univ., Münich, 1971]). Isomorphic copies of each term of the formal deformation are extracted by Grauert’s ideal extension theory (which is briefly sketched). Convergence is proven using this extension theory and a technical cohomological device (Glättungssatz).

I. F. Donin, H. Grauert and P. Deligne have independently obtained the same result. The motivation for this work is therefore the presentation of the Glättungssatz method in place of potential-theoretic arguments, in order to find deformations over singular analytic spaces. The vector bundle case, according to the authors, is particularly simple in relation to the general case of a principal bundle with complex Lie structure group, which will presumably be handled by similar, albeit technically more complicated, methods in a later paper.

Andrew Markoe

MR0361161 (50 13607) 32G05 32G13 Forster, Otto; Knorr, Knut Ein neuer Beweis des Satzes von Kodaira-Nirenberg-Spencer. (German) Math. Z. 139 (1974), 257–291. Kodaira, Nirenberg und Spencer haben 1958 bewiesen, daß zu jeder kompakten komplexen Mannigfaltigkeit X mit H^{2}(X, Θ) = 0 eine verselle Deformation existiert. 1962 hat Kuranishi die Existenz einer versellen Deformation für beliebige kompakte komplexe Mannigfaltigkeiten nachgewiesen; 1974 bewies Grauert eine analoge Aussage für kompakte komplexe Räume.

In der vorliegenden Arbeit wird ein neuer Beweis des Satzes von Kodaira-Nirenberg-Spencer gegeben. Er verwendet Potenzreihenmethoden und Glättungstechniken und kommt ohne Potentialtheorie und fastkomplexe Strukturen aus. Die hier entwickelten Methoden lassen sich auch verwenden, um die Existenz einer versellen Deformation eines kompakten komplexen Raumes X zu beweisen. Die Verfasser stellen den Spezialfall, daß X singularitätenfrei ist und H^{2}(X, Θ) = 0 gilt, hier eigens dar, um die Beweisideen klarer hervortreten zu lassen. Im einzelnen gliedert sich die Arbeit folgendermaßen: § 1 bringt Aussagen über filtriert differenzierbare Abbildungen, § 2 enthält einen Majorantenkalkül für die Komposition von Abbildungen. In § 3 wird ein Theorem B für die Garbe der vertikalen Automorphismen bewiesen; in § 4 ein Glättungssatz. § 5 bringt vor allem die formale Theorie der Deformationen kompakter komplexer Mannigfaltigkeiten: Die Kodaira-Spencer-Abbildung und Aussage über Hindernisse gegen die Erweiterung von Deformationen. In § 6 wird dann die Konstruktion einer versellen Deformation durchgeführt.

H. Kerner

MR0325998 (48 4344) 32C35 Forster, O.; Knorr, K. Relativ-analytische Räume und die Kohärenz von Bildgarben. (German) Invent. Math. 16 (1972), 113–160. In dieser Arbeit wird der Grauertsche Projektionssatz über die Kohärenz der Bildgarben bei eigentlichen holomorphen Abbildungen auf die sogenannten relativ-analytischen Räume verallgemeinert. Hierbei sind als Parameterräume insbesondere auch C^{k}-Mannigfaltigkeiten, 0 ≤ k ≤ ∞, zugelassen. Der Beweisgang für den relativ-analytischen Fall entspricht dem Beweis dem der Verfasser [Manuscripta Math. 5 (1971), 19–44; MR0308437 (46 #7551)] für den anfangs erwähnten Satz von Grauert. Eine ähnliche Darstellung wurde auch von R. Kiehl [Invent. Math. 16 (1972), 40–112] gegeben.

G. Trautmann MR0308438 (46 7552) 32C35 Knorr, Knut Der Grauertsche Projektionssatz. (German) Invent. Math. 12 (1971), 118–172. The author presents a proof of the Grauert projection theorem [H. Grauert, Inst. Hautes Études Sci. Publ. Math. No. 5 (1960); MR0121814 (22 #12544)], following the ideas of Grauert’s original proof. It is the author’s third version of the “classical” proof [see also the author, Ann. Scuola Norm. Sup. Pisa (3) 22 (1968), 729– 761; MR0249666 (40 #2908a); ibid. (3) 23 (1969), 1–74; MR0249667 (40 #2908b)]. Although the main ideas are those of Grauert, there are many improvements, for example a more elegant proof of the theorem on Grauert-privileged polydiscs, a new notion of “Meßüberdeckung” and many other improvements in details. In spite of the new proofs by R. Kiehl and J.-L. Verdier [Math. Ann. 195 (1971), 24– 50; MR0306555 (46 #5678)] and O. Forster and the author [#7551 above], it is worth looking at this elegant and relatively short version of Grauert’s proof.

G. Trautmann

MR0308437 (46 7551) 32C35 Forster, O.; Knorr, K. Ein Beweis des Grauertschen Bildgarbensatzes nach Ideen von B. Malgrange. (German. English summary) Manuscripta Math. 5 (1971), 19–44. The authors give a proof of the Grauert direct image theorem [see H. Grauert, Inst. Hautes Études Sci. Publ. Math. No. 5 (1960); MR0121814 (22 #12544)] that is different from Grauert’s original proof and avoids many technical difficulties. The most important feature of the authors’ proof is that it is not necessary any more to do induction on the dimension of the base space. Also it is not necessary to make use of the theorem of privileged polydiscs. Let π: X → S be a proper holomorphic map of complex spaces, assume without loss of generality that S is open in C^{N}, and let ℱ be a coherent 𝒪_{X}-sheaf on X. In order to prove the coherence of the direct image sheaves R^{n}π_{∗}ℱ, the authors replace the usual Čech-complex of ℱ on parts of X by a larger complex C^{·} = C^{·}(r, U, ℱ), U ⊂ S open, with certain real parameters r, such that C^{·} can be considered as the section complex Γ(U, 𝒞^{·}(r)) of a complex 𝒞^{·}(r) of 𝒪_{S}-sheaves and such that ℋ^{n}𝒞^{·}(r) = R^{n}π_{∗}ℱ for small r. This complex also contains all information about local resolutions of ℱ by free sheaves of finite rank. Moreover, the complex C^{·} is constructed in such a way that the “Meßüberdeckungen” of Grauert and his Leray-lemma with bounds are no longer necessary. The main step in the authors’ proof is the construction of a second complex ℒ^{·} of free 𝒪_{S}-sheaves of finite rank on S that is quasi-isomorphic to 𝒞^{·} so that finally R^{n}π_{∗}ℱ = ℋ^{n}𝒞^{·} = ℋ^{n}ℒ^{·} is 𝒪_{S}-coherent. The construction of ℒ^{·} is done inductively by using norms for analytic entities that are analogous to the original norms of Grauert and by proving implicitly the nuclearity of certain restriction mappings. The construction of ℒ^{·} is equivalent to a similar construction due to R. Kiehl and J. L. Verdier [Math. Ann. 195 (1971), 24–50; MR0306555 (46 #5678)], who make a systematic and extensive use of nuclear mappings.

G. Trautmann

MR0293129 (45 2208) 32F10 Knorr, Knut; Schneider, Michael Relativexzeptionelle analytische Mengen. (German) Math. Ann. 193 (1971), 238–254. The authors generalize H. Grauert’s result on exceptional analytic sets [same Ann. 146 (1962), 331–368; MR0137127 (25 #583)] to the relative case and thereby obtain the relative Kodaira vanishing theorem and a criterion for holomorphic maps to be projective. The proof consists in replacing the finiteness theorem for strictly pseudoconvex spaces used in Grauert’s original proof by the coherence theorem for 1-convex maps, which is derived from a coherence result of Knorr for convex maps and a relative Remmert reduction (compare the reviewer’s paper [ibid. 190 (1970/71), 203–214; MR0279341 (43 #5063)]).

Y.-T. Siu

MR0277764 (43 3497) 32.50 Knorr, K. Le théorème de projection de Grauert. (French) 1971 Espaces Analytiques (Séminaire, Bucharest, 1969) pp. 17–24 Editura Acad. R.S.R., Bucharest The theorem in the title says that the direct image R^{n}π_{∗}ℱ (n ≥ 0) of a coherent analytic sheaf ℱ on X is also coherent if the mapping π: X → S is proper. The author gives a new proof of the above theorem, and some related topics. He begins with some auxiliary results and his proof is essentially a double induction on k and n. Here the parameter k arises from the fact that the support of ℱ is contained in π^{−1}T, where T is a simple analytic set of codimension k in S.

S. Hitotumatu

MR0276497 (43 2244) 32.50 Knorr, K. Die Endlichkeitssätze für q-konvexe und p-konkave Abbildungen und einige Anwendungen. (German. French summary) 1971 Espaces Analytiques (Séminaire, Bucharest, 1969) pp. 9–15 Editura Acad. R.S.R., Bucharest The theorem in the title is due to A. Andreotti and H. Grauert [Bull. Soc. Math. France 90 (1962), 193–259; MR0150342 (27 #343)] and is as follows. For an N-dimensional complex space X, suppose that there exist a function φ: X → R and constants −∞ ≤ a < b < +∞ with the following properties: (1) {φ < a} is strongly p-convex and {φ > b} is strongly q-convex, and (2) for all a′ ≤ a < b ≤ b′ both {a′ ≤ φ ≤ b} and {a ≤ φ ≤ b′} are compact; then for every coherent analytic sheaf ℱ on X, the cohomology H^{n}(X, ℱ) is a finitedimensional vector space for q ≤ n < prof_{{φ<a}}ℱ − p. Here prof_{U}ℱ is the infimum of N − d_{x} for x ∈ U, where d_{x} is the minimum length of the Hilbert resolutions of the type 0 → 𝒪^{p_{d}} → ⋯ → 𝒪^{p_{2}} → 𝒪^{p_{1}} → ℱ → 0 in the neighborhood of x. Then the author gives a conjecture corresponding to a generalization of Grauert’s projection theorem. Let φ satisfy the above condition (1) and π: X → S(open set)⊂ C^{m} be a mapping that is proper in the sets {a′ ≤ φ ≤ b} and {a ≤ φ ≤ b′} for every a′ ≤ a < b ≤ b′. Then, is the direct image R^{n}π_{∗}ℱ of a coherent analytic sheaf ℱ on X coherent for q ≤ n < prof_{{φ<a}}ℱ − m − p? He shows that it is true for two special cases corresponding to b = +∞ and a = −∞, respectively. Finally, he discusses the applications of the above theorem to the exceptional analytic set and projective holomorphic mapping.

S. Hitotumatu